OpenStudy (anonymous):

use he distributive property to help you solve the equation 7 ( w - 2) = -14 w=

OpenStudy (zarkon):

0

OpenStudy (anonymous):

\[ \begin{eqnarray*} 7(w-2)& =& -14 \\ 7w-7\cdot 2& =& -14 \\ \end{eqnarray*}\]Solve

OpenStudy (anonymous):

is it 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Was going to answer, but I'm too late...sorry :